On January 1, 2017, Pronghorn Company issued 10-year, $2,090,000 face value, 6% bonds, at par. Each $1,000 bond is convertible into 14 shares of Pronghorn common stock. Pronghorn’s net income in 2017 was $311,000, and its tax rate was 40%. The company had 101,000 shares of common stock outstanding throughout 2017. None of the bonds were converted in 2017.

(a) Compute diluted earnings per share for 2017. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share $


(b) Compute diluted earnings per share for 2017, assuming the same facts as above, except that $1,010,000 of 6% convertible preferred stock was issued instead of the bonds. Each $100 preferred share is convertible into 5 shares of Pronghorn common stock. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share $
Answer #1

(a) Interest amount after tax = 2,090,000 x 6% x (1 - 0.40)

= $75,240

Diluted shares = 2,090,000/1,000 x 14 = 29,260 shares

DEPS = (311,000 + 75,240)/(101,000 + 29,260)

= 386,240/130,260

= $2.97 per share

(b) Diluted shares = 1,010,000/100 x 5 = 50,500 shares

DEPS = 311,000/(101,000 + 50,500)

= $2.05 per share